OpenStudy (anonymous):

HAPPY NUMBERS--- 1. add the squares of the digits of a whole number to produce a new number. 2. if you repeat this process and end up with the number 1, you have found the happy number. (e.g. 32 is a happy number) 32- 9+4=13 - 1+9= 10- 1 + 0 = 1 - is an arrow that is an example. Find the happy numbers greater than 40....

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

happy numbers

myininaya (myininaya):

So say we have the number 11 (this is a whole number by the way) \[1^2+1^2=1+1=2\] So since we didn't get 1, 11 is not a happy number But they say 32 is \[3^2+2^2=9+4=13\] \[1^2+3^2=1+9=10\] \[1^2+0^2=1+0=1\] So yes indeed 32 is happy! :) ----- So we want to find a happy number greater than 40... lets think for a sec (or let me think for a second)

myininaya (myininaya):

This is a good question...I think I need help... Let me see if I can get some @imranmeah91 @bahrom7893 @jamesj @zarkon @pre-algebra

OpenStudy (anonymous):

ok

myininaya (myininaya):

it doesn't say from a happy number greater than 40 it says find all happy numbers greater than 40 that scares me a bit

myininaya (myininaya):

find*

OpenStudy (bahrom7893):

lol no need to include my name if myin needs help hahah

OpenStudy (bahrom7893):

Let's try 42.. the answer to everything: 4^2+2^2=16+4=20 nah doesn't work

OpenStudy (bahrom7893):

wait all happy numbers? uh oh... other numbers are getting depressed..

OpenStudy (anonymous):

they just use + and = in this

myininaya (myininaya):

So if 32 is happy, then 320,3200,32000,320000, and so on is...

myininaya (myininaya):

@Zed

myininaya (myininaya):

Bahrom also mentioned 1,10,100,1000,10000,10000, and so on is...

OpenStudy (anonymous):

i am sorry it says find the first two happy numbers greater than 40

myininaya (myininaya):

Oh thank god...

OpenStudy (bahrom7893):

LOL that's like *facepalm*....

OpenStudy (bahrom7893):

41: 4^2+1^2=16+1=17 1^2+7^2=1+49=50 5^2=25 4+25=29 4+81=85 64+25=89 64+81=.. do i keep going like this?

myininaya (myininaya):

41 \[4^2+1^2=16+1=17\] \[1^2+7^2=1+49=50\] \[5^2+0^2=25\] \[2^2+5^2=29\] \[2^2+9^2=85\] \[8^2+5^2=89\] \[8^2+9^2=145\] \[1^2+4^2+5^2=42\] \[4^2+2^2=20\] \[2^2+0^2=4 \neq 1\] So 41 is not happy

OpenStudy (bahrom7893):

lol myin im impressed by your patience

myininaya (myininaya):

and 42 is not happy looking at this also

OpenStudy (bahrom7893):

yea i got that

OpenStudy (mani_jha):

Writing a computer program for this would help greatly, although it would be cheating :P Checking for every number is gonna be a daunting task

myininaya (myininaya):

43 \[4^2+3^2=25\] we showed this was not happy when we were showing that 41 wasn't

myininaya (myininaya):

44 \[4^2+4^2=32\] 32 is happy so 44 is

OpenStudy (bahrom7893):

45=16+25=41 unhappy

myininaya (myininaya):

46 \[4^2+6^2=16+36=52\] \[5^2+2^2=29\] We already showed 29 is not happy when we looked at 41 so 46 is not happy

OpenStudy (anonymous):

did u get the file

myininaya (myininaya):

yes we already found one memelee

myininaya (myininaya):

44

myininaya (myininaya):

Now we are looking at 47

OpenStudy (anonymous):

i know thank you i thought if you could see what i am trying to help my child with it would be easier all thiese r crazy

OpenStudy (bahrom7893):

16+49=65 36+25=61 36+1=37 9+49=68 36+64=100 yay HAPPY!

myininaya (myininaya):

yea we are happy now! :)

OpenStudy (bahrom7893):

check my arithmetic btw

myininaya (myininaya):

Thank you guys for coming and helping memelee and me. I was freakin' out for a sec about finding all happy numbers greater than 40.

myininaya (myininaya):

no memelee 47 he just checked 47 and it was happy

OpenStudy (anonymous):

o lol i am so confused you all r the best

OpenStudy (anonymous):

if anyone would like to look at this attachment any help is welcomed so 44 and 47 R HAPPY

OpenStudy (zarkon):

44, 49

myininaya (myininaya):

44,47

OpenStudy (zarkon):

\[\cancel{47}\]

myininaya (myininaya):

so I need to check bahrom's arithmetic lol

myininaya (myininaya):

9+49=58 not 68

OpenStudy (anonymous):

ok can you all see my attachment i need help with alot of it and i dont know if i sent it lol;

OpenStudy (zarkon):

just so you know...according to my new 'happy' program 103456789298765430297364521 is happy ;)

OpenStudy (anonymous):

thank you i still dont know how to show this to my child and can you see my attachment?

OpenStudy (zarkon):

1034567892987654302973645212657823642782587562742523534535245306 is happy...these are fun ;)

OpenStudy (zarkon):

which one are you working on?

OpenStudy (anonymous):

all

OpenStudy (anonymous):

i am trying to help my child but if i dont know it how can i show her

OpenStudy (zarkon):

it looks like you have answers to a couple of them

OpenStudy (zarkon):

by the way... 1404720637424776171015893285076586022511883944493227893407007626053358017096629130428838656664914644591216791171720794830512795459486045059833726538116396895070853996153616905628590023123719241557756032598746616630754153537124256386206122191082028591423913406779794362892355305148462788276820994110962089346622067555874958960623925865096155643308442373668252970647464664574958103406799756720487147207232270715918418384253720418235996188915183594994097160164143780453560562915252718288547951629007629788316036656675693169670540801230860237114576538746238796835257460586564933573664427518014073628649955202214934233356041506519165812 is happy :)

OpenStudy (anonymous):

yes if they are correct i do believe # 29 is roman catholic

OpenStudy (anonymous):

wow thats alot of happy numbers

OpenStudy (zarkon):

looks right

OpenStudy (zarkon):

I wrote a program on my TI-Nspire CX CAS

OpenStudy (anonymous):

no clue what that is

OpenStudy (zarkon):

which problem are you working on now...and don't say all of them.

OpenStudy (anonymous):

24

OpenStudy (zarkon):

vertical angles are congruent.

OpenStudy (anonymous):

is x 150 also

OpenStudy (zarkon):

yes it is 150

OpenStudy (anonymous):

k so 150 so that is exackly the same

OpenStudy (anonymous):

#25 B

OpenStudy (zarkon):

no

OpenStudy (anonymous):

c cause of the top is the same

OpenStudy (zarkon):

no

OpenStudy (anonymous):

has to be A

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

confused cause the shape for the top looks like C

OpenStudy (zarkon):

but the sides are triangles

OpenStudy (anonymous):

ok

OpenStudy (zarkon):

and the base of A is a rectangle

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

#30 is D

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

# 27, 28, 31, and 32

OpenStudy (anonymous):

# 31 is 15?

OpenStudy (zarkon):

16

OpenStudy (anonymous):

can u show me

OpenStudy (zarkon):

solve \[\frac{3}{2}=\frac{w+8}{w}\] for w

OpenStudy (anonymous):

w is 2

OpenStudy (zarkon):

w=16

OpenStudy (anonymous):

3/2 = w+8 over 2 = 16

OpenStudy (zarkon):

\[\frac{3}{2}=\frac{w+8}{w}\] \[\Rightarrow 3w=2(w+8)\] \[\Rightarrow 3w=2w+16\] \[\Rightarrow w=16\]

OpenStudy (anonymous):

k wrote that down so i can show my child thank you